High-pressure liquid chromatography (HPLC) is a method used in chemistry and biochemistry to purify chemical substances. The pressures used in this procedure range from around 500 kilopascals (500,000 Pa) to about 60,000 kPa (60,000,000 Pa). It is often convenient to know the pressure in torr. If an HPLC procedure is running at a pressure of 4.84×108 Pa , what is its running pressure in torr?
Answer #1

1atm = 760torr = 101325Pa

1 Pa = 760/101325torr

Pressure = 4.84×108 Pa = 4.84 ×108 Pa × (760torr/1atm)× (1atm/101325Pa)

= 0.036303 ×108 Torr

= 3.63×106 Torr. (Answer)